OpenStudy (anonymous):

A prism has a right triangle as its base with a hypotenuse that measures 5 meters and legs that measure 3 meters and 4 meters. What is the surface area of this prism if the height of the prism is 6 meters?

OpenStudy (nurali):

Total surface area has : 2 triangles (one at each end) of area (1/2)(3)(4) = 6 So the surface area of those triangles combined is 12 m^2 It has one side of width 3 and height 6. Surface area of that is (3 * 6) = 18 m^2 It has one side of width 4 and height 6. Surface area of that is (4 * 6) = 24 m^2 It has one side of width 5 and height 6. Surface area of that is (5 * 6) = 30 m^2 Total surface area = (12 + 18 + 24 + 30) = 84 m^2

OpenStudy (anonymous):

A)102 m2 B)96 m2 C)78 m2 D)84 m2

OpenStudy (anonymous):

@Nurali thank you !!! :D

OpenStudy (nurali):

Anytime.